# When Does Learning Renormalize? Sufficient Conditions for Power–Law Spectral Dynamics

### v3

@@ A.1 Boundedness of Particular Recurrent Networks
Choosing $K=K_{\delta}$ so that $\rho^{K_{\delta}}\lesssim\delta$ yields (9). Finally, if we define the GRSD blocks at a coarse scale larger than $K_{\delta}$, the residual $\mathcal{R}^{(l)}_{\delta}$ becomes negligible at the GRSD shell scale, and the exact banded form (10) holds at that renormalized resolution. ∎
@@ A.2 Proof of Theorem 1
In addition, assume the two GRSD structural properties already adopted in the main text: (i) intra–shell couplings are antisymmetric, so shell–internal transfers cancel in the shell energy balance; (ii) the renormalizable closure holds, meaning that inter–shell evolution can be expressed (up to a controlled remainder) via boundary fluxes between neighboring shells.
@@
We prove Theorem 1 by decomposing the argument into three steps. Step I establishes the existence of a renormalized shell dynamics under Conditions 13. Steps II and III, which are new compared to the original formulation, use Condition 4 together with gradient–flow covariance to derive the power–law form of the velocity field.
@@
Let $\{S_{\alpha}\}$ be the GRSD logarithmic shells in $\lambda$ (equivalently, equal–width bins in $s=\log\lambda$). Let $E_{\alpha}(t)$ denote the shell energy (shell–aggregated error energy), and let $\varepsilon(\lambda,t)$ be the corresponding shell–averaged energy density (defined by the usual piecewise–constant interpolation over shells). Let $F_{\alpha+\frac{1}{2}}(t)$ denote the net flux across the boundary between $S_{\alpha}$ and $S_{\alpha+1}$, and let $D_{\alpha}(t)$ be the shell–aggregated dissipation term. The GRSD renormalizable closure asserts that the inter–shell dynamics can be written (up to a controlled remainder) in terms of these boundary fluxes.
Step I: From (1)–(3) to shell conservation with boundary fluxes.
@@
By the bandedness condition (10), we may write for each $l$,
$\dot{J}^{(l)}(t)=\sum_{|p-l|\leq K}A_{lp}(t)\,J^{(p)}(t),$ (15)
for some operator coefficients $A_{lp}(t)$ (depending on the learning configuration and the chosen block basis). By (12), these coefficients are uniformly bounded in operator norm; concretely, there exists $C_{A}<\infty$ such that
@@
Taking operator norms and using submultiplicativity,
@@
Now define a distance–profile envelope for each time $t$:
@@
Using (18) and the fact that |pm||p-m| differs from |lm||l-m| by at most $K$, one obtains the envelope inequality
@@
where we set $u_{k}=0$ for $k<0$. This is a finite–range “mixing” inequality on the discrete line $\mathbb{Z}_{\geq 0}$.
Consider the weighted sum $U(t):=\sum_{k\geq 0}w_{k}u_{k}(t)$ with weights $w_{k}:=\rho^{k}$ for some $\rho\in(0,1)$. Multiplying (19) by $w_{k}$ and summing over $k\geq 0$ gives [evans2022partial, gronwall1919note]
U˙(t) $\dot{U}(t)\;\leq\;2C_{A}\Bigl(\sum_{j=-K}^{K}\rho^{-j}\Bigr)\,U(t)\;=\;C_{\rho}\,$U(t), (20)
for a constant $C_{\rho}<\infty$. By Grönwall [gronwall1919note],
$U(t)\leq e^{C_{\rho}t}U(0).$ (21)
At $t=0$, (11) implies $u_{k}(0)\leq\varepsilon_{k}$ for $k\geq 1$, hence $U(0)<\infty$ since $\sum_{k}\varepsilon_{k}<\infty$. Therefore $U(t)<\infty$ for all $t\in[0,T]$, and in particular we may set
$\tilde{\varepsilon}_{k}:=\rho^{-k}\sup_{t\in[0,T]}U(t),$
to obtain a summable bound of the form
@@
This establishes that long–range block correlations remain integrable along the entire path.
(I.b) Boundary–flux form of the shell energy balance. By the GRSD antisymmetry property, all intra–shell transfers cancel in the evolution of $E_{\alpha}(t)$. Thus,
@@
The renormalizable closure asserts that, up to a remainder controlled by the summable tail (22), the inter–shell transfers are supported near shell boundaries and can be written as a telescoping difference of boundary fluxes:
$\frac{d}{dt}E_{\alpha}(t)=F_{\alpha-\frac{1}{2}}(t)-F_{\alpha+\frac{1}{2}}(t)-D_{\alpha}(t),$ (24)
@@
where $F_{\alpha+\frac{1}{2}}(t)$ depends only on spectral content in a fixed neighborhood of the boundary between $S_{\alpha}$ and $S_{\alpha+1}$. Passing from the discrete balance (24) to a shell–averaged density representation yields the one–dimensional conservative form
where $J(\lambda,t)$ is the flux density consistent with the boundary fluxes $F_{\alpha+\frac{1}{2}}(t)$ under the GRSD log–shell interpolation. This completes Step I. The reduction from inter-shell transfers to boundary fluxes follows standard renormalizable closure arguments in large-scale interacting systems [kolmogorov1995turbulence, forster1977large, spohn2012large].
Step II: Scale covariance as a rigidity consequence of renormalizable closure.
@@
We now invoke Condition 4 (log–shift invariance). On the intermediate spectral window, the renormalized shell coupling statistics entering the GRSD closure are asymptotically translation invariant in $s=\log\lambda$. Equivalently, for any $\tau\in\mathbb{R}$, a shift $\varepsilon(s,t)\mapsto\varepsilon(s+\tau,t)$ induces a corresponding shift of the flux $J(s,t)$, without introducing any intrinsic length scale.
@@
As a consequence, the velocity field $v(s,t)$ cannot depend on the absolute position $s$ except through multiplicative rescaling. In particular, any admissible dependence of $v$ on $s$ must be compatible with the functional relation
@@ A.3.2 Markov–additive structure of log–spectral increments
The pair $(u_{\ell},\delta_{\ell})$ defines a Markov–additive process [meyn2012markov] on $\mathbb{S}^{n-1}\times\mathbb{R}$. Under Condition 3 and the residual small–step assumption, $\delta_{\ell}$ admits uniform moment bounds
@@ A.3.3 Self–averaging of accumulated log–increments
Since $u_{\ell}$ is a uniformly ergodic Markov chain (see Section A.3.4), standard Berry–Esseen bounds [bolthausen1982berryesseen] for Markov–additive processes apply. In particular, there exists $C_{\mathrm{BE}}<\infty$ such that
@@ A.3.4 Directional mixing
For products of i.i.d. random matrices, classical results of Furstenberg and Kesten establish the existence of a unique stationary distribution on the projective space and ergodicity of the induced Markov dynamics, leading to direction independent asymptotic behavior of matrix products under suitable conditions [furstenberg1960products]. The directional updates induced by the random residual perturbations define a Markov process on the unit sphere, whose mixing properties can be analyzed in the classical Markov chain framework [levin2009markov]. Standard coupling or diffusion–approximation arguments imply that the mixing time $\tau_{\mathrm{mix}}(\eta)$ satisfies
@@ A.3.5 Log–bin averaging and kernel emergence
Let $X_{uv}:=\phi_{u}^{\ast}A\phi_{v}$ denote generic quadratic couplings entering the GRSD closure. By Condition 3, conditional second moments satisfy [koltchinskii2017normal]

### v4

@@ A.1 Boundedness of Particular Recurrent Networks
Choosing $K=K_{\delta}$ so that $\rho^{K_{\delta}}\lesssim\delta$ yields (9). Finally, if we define the GRSD blocks at a coarse scale larger than $K_{\delta}$, the residual $\mathcal{R}^{(l)}_{\delta}$ becomes negligible at the GRSD shell scale, and the exact banded form holds at that renormalized resolution. ∎
@@ A.2 Proof of Theorem 1
In addition, assume the GRSD structural property already adopted in the main text: (i) intra–shell couplings are antisymmetric, so shell–internal transfers cancel in the shell energy balance.
@@
We prove Theorem 1 by decomposing the argument into three steps. Step I establishes an effective one–dimensional shell conservation law with nearest–neighbor boundary fluxes, where the closure is obtained as a structural consequence of $M=JJ^{*}$ via a Taylor expansion. Steps II and III use Condition 4 together with gradient–flow covariance to derive the power–law constraint on the velocity field.
@@
Let $\{S_{\alpha}\}$ be the GRSD logarithmic shells in $\lambda$ (equivalently, equal–width bins in $s=\log\lambda$). Let $E_{\alpha}(t)$ denote the shell energy (shell–aggregated error energy), and let $\varepsilon(\lambda,t)$ be the corresponding shell–averaged energy density (defined by the usual piecewise–constant interpolation over shells). Let $F_{\alpha+\frac{1}{2}}(t)$ denote the net flux across the boundary between $S_{\alpha}$ and $S_{\alpha+1}$, and let $D_{\alpha}(t)$ be the shell–aggregated dissipation term.
Step I: From (1)–(3) to shell conservation with nearest–neighbor boundary fluxes.
@@
By (10), for each $l$ there exist operator coefficients $A_{lp}(t)$ such that
$\dot{J}^{(l)}(t)=\sum_{|p-l|\leq K}A_{lp}(t)\,J^{(p)}(t).$ (15)
By (12), these coefficients are uniformly bounded: there exists $C_{A}<\infty$ such that
@@
Taking operator norms and using submultiplicativity gives
@@
Define the distance–profile envelope
@@
Using (18) and that |pm||p-m| differs from |lm||l-m| by at most $K$, one obtains
@@
with $u_{k}=0$ for $k<0$. Let $U(t):=\sum_{k\geq 0}\rho^{k}u_{k}(t)$ for some $\rho\in(0,1)$. Then (19) implies a Grönwall inequality
$\dot{U}(t)\;\leq\;C_{\rho}\,U(t),\qquad C_{\rho}:=2C_{A}\Bigl(\sum_{j=-K}^{K}\rho^{-j}\Bigr)<\infty,$ (20)
hence $U(t)\leq e^{C_{\rho}t}U(0)$. By (11), $U(0)<\infty$ since $\sum_{k}\varepsilon_{k}<\infty$. Therefore $U(t)<\infty$ for all $t\in[0,T]$, and in particular there exists a summable tail $\{\tilde{\varepsilon}_{k}\}_{k\geq 1}$ such that
@@
(I.b) Automatic nearest–neighbor closure via a parameter–space Taylor expansion and the exact $\Omega$ formula. By antisymmetry, intra–shell transfers cancel in the energy balance, hence
@@
We now show that direct transfers between non–adjacent shells are second order, and therefore the inter–shell dynamics closes (up to a controlled remainder) on nearest–neighbor boundary fluxes [^2] The idea is inspired by [44]. The argument uses a Taylor expansion of the model output in parameter space and the exact formula for the spectral rotation generator $\Omega$.
Taylor expansion of the output and the order of $\Delta M$.
Let $f(\theta)$ denote the model output in function space, with Jacobian $J(\theta)=\partial_{\theta}f(\theta)$ and squared loss $\mathcal{L}(\theta)=\tfrac{1}{2}\|e(\theta)\|^{2}$ where $e(\theta)=f(\theta)-y$. For a small parameter increment $\delta\theta$, the output admits the expansion
$f(\theta+\delta\theta)=f(\theta)+J(\theta)\,\delta\theta+\frac{1}{2}\,\mathcal{H}(\theta)[\delta\theta,\delta\theta]+O(\|\delta\theta\|^{3}),$ (23)
where $\mathcal{H}(\theta)$ is the second derivative (a Hessian tensor) of $f$.
Consider one infinitesimal gradient–flow step $\delta\theta=\dot{\theta}(t)\,\delta t$ with $\dot{\theta}(t)=-\nabla_{\theta}\mathcal{L}(\theta(t))=-J(\theta(t))^{*}e(t)$. The linearized (first–order) model $f_{\mathrm{lin}}(\theta+\delta\theta):=f(\theta)+J(\theta)\delta\theta$ keeps $J(\theta)$ fixed, hence the corresponding operator
$M_{\mathrm{lin}}:=J(\theta)J(\theta)^{*}$
is constant over the step. Therefore any change of $M(t)=J(t)J(t)^{*}$ must come from the quadratic and higher terms in (23), i.e. from curvature. Equivalently, over a time increment $\delta t$,
$\Delta M(t):=M(t+\delta t)-M(t)=O(\|\delta\theta\|)=O(\delta t),$ (24)
and the contribution of $\Delta M$ to the error update appears only at second order in time, because the first–order error update uses $M(t)$ frozen:
$e(t+\delta t)=e(t)-M(t)e(t)\,\delta t\;+\;\frac{(\delta t)^{2}}{2}\Bigl(M(t)^{2}e(t)-\dot{M}(t)e(t)\Bigr)\;+\;O((\delta t)^{3}).$ (25)
In particular, the term involving $\dot{M}$ (hence curvature) enters the error evolution only through the $(\delta t)^{2}$ term.
Exact $\Omega$ formula and gap suppression across non–adjacent shells.
Let $\{(\lambda_{i}(t),u_{i}(t))\}$ be an orthonormal eigen-decomposition of $M(t)$ on the GRSD window and define the instantaneous rotation generator
$\Omega_{ij}(t):=\langle u_{i}(t),\dot{u}_{j}(t)\rangle,\qquad\Omega_{ij}=-\Omega_{ji}.$
For $i\neq j$, the standard exact identity gives
$\Omega_{ij}(t)=\frac{\langle u_{i}(t),\dot{M}(t)u_{j}(t)\rangle}{\lambda_{j}(t)-\lambda_{i}(t)}.$ (26)
Assume the GRSD shell partition $\{S_{\alpha}\}$ is chosen so that there exists a minimal spectral gap between non–adjacent shells:
$\inf_{t\in[0,T]}\ \inf_{\begin{subarray}{c}i\in S_{\alpha},\,j\in S_{\beta}\\ |\alpha-\beta|\geq 2\end{subarray}}|\lambda_{j}(t)-\lambda_{i}(t)|\geq\Delta.$ (27)
Under the controlled path condition (12) we have $\sup_{t\in[0,T]}\|\dot{M}(t)\|_{\mathrm{op}}<\infty$, hence (26) and (27) imply the uniform bound
$\sup_{t\in[0,T]}\ \sup_{\begin{subarray}{c}i\in S_{\alpha},\,j\in S_{\beta}\\ |\alpha-\beta|\geq 2\end{subarray}}|\Omega_{ij}(t)|\ \leq\ \frac{\sup_{t\in[0,T]}\|\dot{M}(t)\|_{\mathrm{op}}}{\Delta}.$ (28)
Second–order non–adjacent transfers and boundary–flux closure.
Write the error in the instantaneous eigenbasis, $e(t)=\sum_{i}a_{i}(t)u_{i}(t)$, and note that inter–shell exchange is induced by basis rotation. Combining the time–Taylor expansion (25) with the fact that $\Omega$ is generated by $\dot{M}$ through (26), we obtain the key structural consequence: since $\dot{M}$ enters $e(t+\delta t)$ only at order $(\delta t)^{2}$, any energy exchange mediated by $\Omega$ is at least second order in $\delta t$. Moreover, for non–adjacent shells $|\alpha-\beta|\geq 2$, the gap bound (27) controls the corresponding rotation coefficients via (28), so the associated transfers are uniformly of second order:
$T_{\alpha\beta}(t)=O((\delta t)^{2})\cdot O\!\Bigl(\Delta^{-1}\sup_{t\in[0,T]}\|\dot{M}(t)\|_{\mathrm{op}}\Bigr),\qquad|\alpha-\beta|\geq 2,$ (29)
with constants controlled by the bounds in Conditions (1)–(3). (Adjacent shells are not covered by the uniform gap bound and therefore may contribute at leading order; these contributions define the renormalized boundary fluxes.)
Consequently, the inter–shell balance (22) closes, up to a controlled remainder collecting the non–adjacent second–order transfers, in telescoping boundary–flux form
$\frac{d}{dt}E_{\alpha}(t)=F_{\alpha-\frac{1}{2}}(t)-F_{\alpha+\frac{1}{2}}(t)-D_{\alpha}(t)+R_{\alpha}(t),$ (30)
@@
where $F_{\alpha+\frac{1}{2}}(t)$ depends only on spectral content in a fixed neighborhood of the boundary between $S_{\alpha}$ and $S_{\alpha+1}$, and the remainder $R_{\alpha}(t):=\sum_{|\beta-\alpha|\geq 2}T_{\alpha\beta}(t)$ is uniformly controlled on $[0,T]$ by the second–order estimate (29) together with the summable incoherence propagation from Step I.a. Passing to the shell–averaged density representation yields the effective conservation law
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}J(\lambda,t)=-D(\lambda,t)+r(\lambda,t),$ (31)
where $r$ corresponds to the controlled remainder. This completes Step I.
Step II: Scale covariance as a rigidity consequence of log–shift invariance.
@@
We now invoke Condition 4 (log–shift invariance). On the intermediate spectral window, the renormalized shell coupling statistics governing the effective flux law (31) are asymptotically translation invariant in $s=\log\lambda$. Equivalently, for any $\tau\in\mathbb{R}$, a shift $\varepsilon(s,t)\mapsto\varepsilon(s+\tau,t)$ induces a corresponding shift of the flux $J(s,t)$, without introducing any intrinsic length scale.
@@
As a consequence, the velocity field $v(s,t):=J(s,t)/\varepsilon(s,t)$ cannot depend on the absolute position $s$ except through multiplicative rescaling. In particular, any admissible dependence of $v$ on $s$ must be compatible with a functional relation of the form
@@ A.3.2 Markov–additive structure of log–spectral increments
The pair $(u_{\ell},\delta_{\ell})$ defines a Markov–additive process [33] on $\mathbb{S}^{n-1}\times\mathbb{R}$. Under Condition 3 and the residual small–step assumption, $\delta_{\ell}$ admits uniform moment bounds
@@ A.3.3 Self–averaging of accumulated log–increments
Since $u_{\ell}$ is a uniformly ergodic Markov chain (see Section A.3.4), standard Berry–Esseen bounds [3] for Markov–additive processes apply. In particular, there exists $C_{\mathrm{BE}}<\infty$ such that
@@ A.3.4 Directional mixing
For products of i.i.d. random matrices, classical results of Furstenberg and Kesten establish the existence of a unique stationary distribution on the projective space and ergodicity of the induced Markov dynamics, leading to direction independent asymptotic behavior of matrix products under suitable conditions [8]. The directional updates induced by the random residual perturbations define a Markov process on the unit sphere, whose mixing properties can be analyzed in the classical Markov chain framework [31]. Standard coupling or diffusion–approximation arguments imply that the mixing time $\tau_{\mathrm{mix}}(\eta)$ satisfies
@@ A.3.5 Log–bin averaging and kernel emergence
Let $X_{uv}:=\phi_{u}^{\ast}A\phi_{v}$ denote generic quadratic couplings entering the GRSD closure. By Condition 3, conditional second moments satisfy [27]
@@ A.4 Proof of Proposition 1
Appendix B Spectral operator estimation via block Lanczos-SLQ
B.1 Motivation
Directly forming the operator
$M(\theta)=\mathbb{E}_{x}\big[J(x;\theta)J(x;\theta)^{\top}\big]$
is infeasible for modern neural networks due to its dimensionality. However, our analysis only requires access to spectral information and low-rank projections of $M$ and its directional derivative.
The block Lanczos method provides an efficient way to approximate the action of $M$ on a low-dimensional Krylov subspace using only matrix–vector products. This makes it particularly well-suited for probing spectral structure in high-dimensional learning systems.
B.2 Matrix–vector products via automatic differentiation
For any vector $v$ in error space, the product $Mv$ can be written as
$Mv=\mathbb{E}_{x}\big[J(x)(J(x)^{\top}v)\big].$
This can be evaluated without explicitly forming $J$ by composing a vector–Jacobian product (VJP) with a Jacobian–vector product (JVP).
In practice, for each mini-batch we compute
$u=J(x)^{\top}v,\qquad w=J(x)u,$
using automatic differentiation. A small number of fixed mini-batches is sufficient to obtain a stable Monte Carlo estimate of $Mv$ for the purposes of spectral diagnostics.
B.3 Block Lanczos projection
Algorithm 1 Block Lanczos spectral projection
1:Matrix–vector product oracle $v\mapsto Mv$, block size $b$, steps $k$
2:Initialize block $V_{0}\in\mathbb{R}^{d\times b}$ and orthonormalize
3:for $j=1$ to $k$ do
4:  $W\leftarrow MV_{j-1}$
5:  Orthogonalize $W$ against $\{V_{0},\dots,V_{j-1}\}$
6:  QR factorization $W=V_{j}R_{j}$
7:end for
8:Assemble block tridiagonal matrix $T$
9:Compute eigenpairs $(\lambda_{i},u_{i})$ of $T$
B.4 Ritz basis and projected dynamics
The eigenvectors of the block tridiagonal matrix $T$ define a Ritz basis that approximates the true eigenmodes of $M$ within the Krylov subspace. Expressing operators in this basis yields a meaningful discretization of spectral dynamics while remaining computationally tractable.
All subsequent diagnostics, including shell binning and Toeplitz analysis, are performed in this approximate spectral basis.
B.5 Projected operator derivative
To probe spectral transport, we compute a finite-difference approximation of the directional derivative of $M$ along the training trajectory:
$\dot{M}\approx\frac{M(\theta-\varepsilon\nabla\mathcal{L})-M(\theta)}{\varepsilon}.$
Projecting $\dot{M}$ into the Ritz basis yields the operator
$\Omega=U^{\top}(Q^{\top}\dot{M}Q)U,$
which captures how spectral modes exchange energy under training. This operator forms the basis of all shell-level coupling measurements.
B.6 Shell binning and Toeplitz diagnostics
Eigenvalues are grouped into logarithmic spectral shells. The projected operator $\Omega$ is coarse-grained into a shell coupling matrix $K$ by averaging squared couplings within each shell pair.
Toeplitz structure is assessed by comparing $K$ to its best Toeplitz approximation, obtained by averaging along shell offsets. This diagnostic directly tests the approximate shift-invariance across shells assumed in Condition 4.